# Few-Shot Learning on Graphs: from Meta-learning to Pre-training and Prompting

### v3

@@ 3.1 Strategies for the pre-training stage
Generative strategies. Concurrent to the development of contrastive strategies, generative methods present a different perspective for graph pre-training. These methods aim to reconstruct parts of the graph, unlike contrastive methods that rely on discriminating different instances on the graph. Various reconstruction objectives have been proposed, including the adjacency matrix Kipf and Welling (2016), masked edges Hu et al. (2020b) or node features Hou et al. (2022). These objectives involve structure reconstruction or feature reconstruction, with many studies employing objectives from both categories, as summarized in Table 3.
@@
Structure reconstruction focuses on the graph topology, aiming to regenerate the entire graph structure or part of it, compelling the graph encoder to preserve the inherent structural patterns. For instance, VGAE Kipf and Welling (2016) attempt to reconstruct the adjacency matrix using the inner product of node embeddings. However, it mainly captures one-hop connectivity patterns, overlooking complex high-order structural information. Alternatively, GPT-GNN Hu et al. (2020b) choose to mask a portion of the edges, and aim to reconstruct these missing edges using the remaining unmasked graph structure. Notably, GPT-GNN pursues a dual goal, which includes the generation of node features. The additional objective helps the encoder recognize the dependency between features and structures, potentially facilitating the modeling of higher-order graph structures.
@@
Feature reconstruction shifts the focus to generating node features in either the original feature space or the latent space. LaGraph Xie et al. (2022) introduces a predictive framework to perform unobserved latent graph prediction, employing node masking to compute the invariance term in its predictive objective. Meanwhile, GraphMAE Hou et al. (2022) turns to masked node feature reconstruction, proposing a masking strategy and error metric designed for robust training. Building upon GraphMAE, GraphMAE2 Hou et al. (2023) introduces two decoding strategies, namely, multi-view random re-mask decoding and latent representation prediction, to regularize feature reconstruction and minimize overfitting.
3.2 Downstream adaptation by fine-tuning
@@
The pre-trained graph encoder entails prior knowledge about the intrinsic properties of the graphs used in the pre-training stage. This prior can be transferred to various downstream tasks by initializing a downstream model with the pre-trained weights. To tailor the initial model to each specific task, it is equipped with a task-specific projection head, and undergoes an adaptation stage known as fine-tuning. During fine-tuning, for each task, the model is further trained with task-specific annotations, updating all parameters in both the pre-trained graph encoder and the task head in the process Qiu et al. (2020); Hu et al. (2020b), as shown in Fig. 3(c), left.
@@
However, adaptation by fine-tuning, while intuitive and straightforward, has its limitations Liu et al. (2023a). First, an objective gap often exists between the pretext tasks in pre-training and the actual downstream tasks, which can compromise the effectiveness of the adaptation process. Second, updating all parameters can be inefficient, especially in the case of large pre-trained models. In our context of few-shot learning, fine-tuning may lead to overfitting due to the limited number of task-specific labels Liu et al. (2023c).
3.3 Parameter-efficient downstream adaptation
@@
GPPT Sun et al. (2022) token-subgraph similarity calculation: $\text{sim}(\mathbf{s}_{v},\vec{t}_{y})$ input $\times$ $\times$ random $\checkmark$ $\times$ $\times$
@@
GraphPrompt Liu et al. (2023c) subgraph similarity calculation: $\text{sim}(\mathbf{s}_{u},\mathbf{s}_{v})$ readout $\times$ $\times$ random $\checkmark$ $\checkmark$ $\checkmark$
GraphPrompt+ Yu et al. (2023b) subgraph similarity calculation: $\text{sim}(\mathbf{s}_{u},\mathbf{s}_{v})$ all layers $\times$ $\times$ random $\checkmark$ $\checkmark$ $\checkmark$
GPF Fang et al. (2023) universal feature space (implicit template) input $\times$ $\times$ random $\checkmark$ $\checkmark$ $\checkmark$
@@
MultiGPrompt Yu et al. (2023c) node similarity calculation: $\text{sim}(\mathbf{h}_{u},\mathbf{h}_{v})$ all layers $\times$ $\checkmark$ pre-trained tokens $\checkmark$ $\checkmark$ $\checkmark$
@@
SGL Zhu et al. (2023) dual-template: $\text{CL}(\mathbf{h}_{u},\mathbf{h}_{v}),\text{GL}(x_{v},\tilde{x}_{v})$$\times$ $\checkmark$ $\checkmark$ random $\times$ $\times$ $\checkmark$
SAP Ge et al. (2023) dual-view similarity: $\text{sim}(\text{MLP}(X),\text{GNN}(X,A))$$\times$ $\checkmark$ $\checkmark$ random $\checkmark$ $\times$ $\checkmark$
ULTRA-DP Chen et al. (2023) node-node / node-group similarity calculation: $\text{sim}(\mathbf{h}_{u},\mathbf{h}_{v})$ input $\checkmark$ $\checkmark$ random $\checkmark$ $\checkmark$ $\times$
@@
HGPrompt Yu et al. (2023a) task template: $\text{sim}(\mathbf{s}_{u},\mathbf{s}_{v})$ graph template: $HG\to\{G^{0},G^{1},G^{2},\ldots\}$ readout $\times$ $\times$ random $\checkmark$ $\checkmark$ $\checkmark$
@@
VNT Tan et al. (2023) node attribute reconstruction: $\text{MSE}(x_{v},\tilde{x}_{v})$ structure recovery: $\text{MSE}(\{\mathbf{h}_{u},\mathbf{h}_{v}\})$ input $\times$ $\times$ meta-trained $\checkmark$ $\times$ $\times$
@@
ProG Sun et al. (2023) subgraph classification: $\text{CLS}(\mathbf{s})$$\times$ $\checkmark$ $\times$ meta-trained $\checkmark$ $\checkmark$ $\checkmark$
Table 4: Summary of prompt tuning on graphs. For ease of comparison, we adapt the notations used in the original references with a consistent set of symbols. $v,u$: Nodes $v,u$; $\vec{t}_{y}$: Prototype token $\mathbf{t}$ of class $y$; $\vec{h}_{v}$: Node $v$’s embedding; $x_{v}$: Node $v$’s attribute; $\mathbf{s}_{v}$: Subgraph of $v$’s embedding; CL: Contrastive learning; GL: Generative learning; CLS: Classification; $G$: Homogeneous graph; $HG$: Heterogeneous graph.
@@
Prompting on graphs. Originating in the field of natural language processing (NLP), prompting has demonstrated its efficacy in adapting pre-trained language models to diverse language tasks Liu et al. (2023a). In essence, a prompt vector is introduced to modify or reformulate the original input for the pre-trained model, drawing the downstream task closer to the pretext tasks under a unified template. Notably, the prompt vector has a negligible size relative to the pre-trained parameters. During the adaptation process, only a very small prompt vector is tuned without updating the pre-trained model, making the approach parameter efficient. Recently, prompt tuning have been extended to graph learning Sun et al. (2022); Liu et al. (2023c). Building on a unified template that aligns the pretext and downstream losses, these methods design prompts to further narrow the gap between pretext and downstream tasks, as illustrated in Fig. 3(c), right. Specific approaches vary in the choice of the template and prompt design, as summarized in Table 4.
@@
In terms of template, similarity calculation has been widely adopted Sun et al. (2022); Liu et al. (2023c); Yu et al. (2023b, c); Chen et al. (2023); Yu et al. (2023a) to unify three major tasks: node classification (NC), link prediction (LP), and graph classification (GC). For instance, GraphPrompt Liu et al. (2023c) reformulates LP as calculating the similarity between the ego subgraphs of two nodes. Then, it applies the same template to NC and GC, treating them as calculating the similarity between the subgraph of the target instance (using an ego network for NC, or the target graph itself for GC) and the prototypical subgraph for each class. While GraphPrompt only adopts LP as the pretext task, GraphPrompt+ Yu et al. (2023b) illustrates that other prevailing contrastive pretext tasks can also be redefined using subgraph similarity. Likewise, SGL Zhu et al. (2023) and SAP Ge et al. (2023) also employ a form of similarity calculation for contrastive tasks. SGL further incorporates a generative pretext task based on mask node feature reconstruction, which is employed as an additional template for downstream prompt learning. In contrast, GPF Fang et al. (2023) does not employ an explicit template to align task formats. Instead, it assumes a universal feature space as an implicit template to unify different tasks.
@@
In terms of prompt design, it can be further categorized into feature prompt and structure prompt. On one hand, similar to prompt in NLP Liu et al. (2023a), feature prompt modifies node features, reformulating the original downstream tasks to narrow the gap from the pre-trained model. Some studies incorporate feature prompts in the input layer Sun et al. (2022); Fang et al. (2023); Chen et al. (2023) to explicit modify the input in a task-specific manner, while others apply them to the readout layer to guide subgraph readout Liu et al. (2023c); Yu et al. (2023a), or all layers of the pre-trained model Yu et al. (2023b, c). On the other hand, a structure prompt acts as a virtual graph structure to reformulate the implicit structural input for downstream tasks. Structure prompts can take the form of virtual nodes or graphs representing class prototypes for classification tasks Zhu et al. (2023); Ge et al. (2023); Chen et al. (2023), or virtual nodes that augment link prediction Chen et al. (2023).
To learn a comprehensive range of prior knowledge from diverse perspectives, some studies propose to utilize multiple pretext tasks Zhu et al. (2023); Ge et al. (2023); Chen et al. (2023). However, directly combining various pretext tasks may result in task interference Yu et al. (2023c), compromising the complementary nature of employing multiple tasks. To address this issue, MultiGPrompt Yu et al. (2023c) utilizes a series of pretext tokens to facilitate cooperation among pretext tasks. It further proposes a composed prompt, which is initialized based on the pre-trained pretext tokens, aiming to extract fine-grained knowledge specific to each pretext task from the pre-trained model.
Finally, prompting has also been considered on semantic-rich graphs, including heterogeneous and text-attributed graphs. HGPrompt Yu et al. (2023a) extends GraphPrompt to heterogeneous graphs, utilizing an additional graph template to convert a heterogeneous graph into homogeneous ones, bridging the gap between the two kinds of graph. For text-attributed graphs, G2P2 Wen and Fang (2023a) integrates graph neural networks with a language model, adopting a multi-modal alignment framework Radford et al. (2021) to pre-train the language model and graph model jointly. Alternatively, OFA Liu et al. (2024) describes subgraphs using texts, which can be subsequently fed to a pre-trained language model. Both works employ language-based templates and prompts akin to those used in NLP.

### v4

@@ V-A2 Generative Strategies
Concurrent to the development of contrastive strategies, generative methods present a different perspective for graph pre-training. These methods aim to reconstruct parts of the graph, unlike contrastive methods that rely on discriminating different instances on the graph. Various reconstruction objectives have been proposed, including the adjacency matrix [43], masked edges [39] or node features [163]. These objectives involve structure reconstruction or feature reconstruction, with many studies employing objectives from both categories, as summarized in Table VII. Additionally, similar to contrastive methods, diverse generative methods have also been proposed for various types of graph.
@@
Structure reconstruction focuses on the graph topology, aiming to regenerate the entire graph structure or part of it, compelling the graph encoder to preserve the inherent structural patterns. For instance, VGAE [43] attempt to reconstruct the adjacency matrix using the inner product of node embeddings. However, it mainly captures one-hop connectivity patterns, overlooking complex high-order structural information. Alternatively, GPT-GNN [39] choose to mask a portion of the edges, and aim to reconstruct these missing edges using the remaining unmasked graph structure. Similarly, MaskGAE [77] masks a subset of edges and aims to reconstruct both the masked edges and the degrees of the associated nodes. Beyond reconstructing individual edges, NWR-GAE [161] reconstructs the entire neighborhood of a node, including node degrees and the distribution of neighbors’ features, offering a more comprehensive insight into the graph structure.
@@
Feature reconstruction shifts the focus to generating node features in the original or latent space. LaGraph [162] introduces a predictive framework to perform unobserved latent graph prediction, employing node masking to compute the invariance term in its predictive objective. Meanwhile, GraphMAE [163] turns to masked node feature reconstruction, proposing a masking strategy and error metric designed for robust training. GraphMAE2 [78] further introduces two decoding strategies, namely, multi-view random re-mask decoding and latent representation prediction, to regularize feature reconstruction and minimize overfitting.
Graph types. Different generative pre-training strategies have been developed for diverse types of graph. For knowledge graphs (KGs), Liu et al. [164] tackle complex logical queries by framing it as a masked entity prediction problem. It randomly samples subgraphs from a knowledge graph and masks certain entities for prediction. Besides masking entities, Wen et al. [79] mask some relations to derive a more comprehensive understanding of semantic knowledge. Moreover, MPKG [165] integrates four pretext tasks: knowledge reconstruction, higher-order neighbor reconstruction, feature reconstruction, and relation adaptation to capture heterogeneous item semantics, particularly enhancing performance in zero-shot scenarios. For dynamic graphs, PT-DGNN [166] first samples subgraphs in a time-aware manner as the input to pre-training, and then predicts unobserved edges using observed one to capture the evolutionary traits of the network. On the other hand, STEP [167] masks node features over a time period for prediction, providing an alternative approach to understanding graph evolution. For multi-modal graphs (MMGs), PMGT [168] employs graph structure reconstruction and masked node feature reconstruction, integrating item relations and multi-modal side information. Similarly, ColdGPT [169] designs multiple pretext tasks from various data sources, such as item content, historical purchase sequences, and review texts.
V-A3 Integration with Pre-trained Large Models
Large language models (LLMs) have recently emerged as a powerful and widely adopted paradigm across the broader fields of machine learning and artificial intelligence due to their emergent capabilities. Inspired by their remarkable success, initial efforts have been made to incorporate LLMs into graph pre-training. These explorations predominantly focus on text-attributed graphs (TAGs), where each node or edge is accompanied by textual descriptions that can be naturally processed by LLMs. To integrate both structural and semantic information within TAGs, contrastive pre-training methods have been adopted in several studies [44, 82, 170, 171]. Similar to Contrastive Language-Image Pretraining (CLIP) [172], their core principle is to align text embeddings, derived from a language model, with their corresponding node embeddings produced by a graph encoder. Alternatively, Patton [173] combines masked language modeling [37] and masked node prediction simultaneously to pre-train the LLM and the graph encoder. LLMs can also be used at various stages of graph pre-training. For instance, both GALM [174] and InstructGLM [175] employ link prediction as a pretext task, but their application of LLMs differ. Specifically, GALM feeds LLM-encoded node embeddings into the graph encoder to incorporate graph structural information, using link prediction to pre-train both the LLM and the graph encoder. In contrast, InstructGLM relies on natural language to describe the graph structures and employs the LLM to encode node embeddings. Moreover, instead of relying on unsupervised pre-training strategies, SimTeG [176] first fine-tunes a pre-trained language model with supervision from a downstream task, and then utilizes the fine-tuned model to obtain node embeddings, which can be subsequently used by any graph encoder for training on the same task.
V-B Downstream Adaptation by Fine-tuning
@@
The pre-trained graph encoder entails prior knowledge about the intrinsic properties of the graphs used in the pre-training stage. This prior can be transferred to various downstream tasks by initializing a downstream model with the pre-trained weights. To tailor the initial model to each specific task, it is equipped with a task-specific projection head, and undergoes an adaptation stage known as fine-tuning. During fine-tuning, for each task, the model is further trained with task-specific annotations, which can update the parameters in both the pre-trained model and the task head in the process [30, 39], as shown in Fig. 6(c), left.
@@
However, adaptation by fine-tuning, while intuitive and straightforward, has its limitations [177]. First, an objective gap often exists between the pretext tasks in pre-training and the actual downstream tasks, which can compromise the effectiveness of the adaptation process. Second, updating all parameters can be inefficient, especially in the case of large pre-trained models. In our context of few-shot learning, fine-tuning may lead to overfitting due to the limited number of task-specific labels [19].
V-C Parameter-efficient Downstream Adaptation
@@
TABLE VIII: Summary of prompt tuning on graphs. For ease of comparison, we adapt the notations used in the original references with a consistent set of symbols. $v,u$: Nodes $v,u$; $\vec{t}_{y}$: Prototype token $\mathbf{t}$ of class $y$; $\vec{h}_{v}$: Node $v$’s embedding; $\mathbf{x}_{v}$: Node $v$’s attribute; $\mathbf{s}_{v}$: Subgraph of $v$’s embedding; CL: Contrastive learning; GL: Generative learning; CLS: Classification.
@@
GPPT [48] subgraph-token similarity: $\text{sim}(\mathbf{s}_{v},\vec{t}_{y})$ input $\times$ $\times$ random $\checkmark$ $\times$ $\times$
VPGNN [178] node-token matching: $\text{match}(\mathbf{h}_{v},\mathbf{t}_{y})$$\times$ $\checkmark$ $\times$ random $\checkmark$ $\times$ $\times$
@@
GraphPrompt [19] subgraph similarity: $\text{sim}(\mathbf{s}_{u},\mathbf{s}_{v})$ readout $\times$ $\times$ random $\checkmark$ $\checkmark$ $\checkmark$
MOP [179] readout $\times$ $\times$ random $\times$ $\checkmark$ $\times$
GraphPrompt+ [80] all layers $\times$ $\times$ random $\checkmark$ $\checkmark$ $\checkmark$
ProNoG [180] readout $\times$ $\times$ conditional $\checkmark$ $\checkmark$ $\checkmark$
@@
MDGPT [181] node similarity: $\text{sim}(\mathbf{h}_{u},\mathbf{h}_{v})$ readout $\times$ $\times$ pretext tokens $\checkmark$ $\checkmark$ $\checkmark$
MultiGPrompt [84] all layers $\times$ $\checkmark$ pretext tokens $\checkmark$ $\checkmark$ $\checkmark$
HetGPT [116] input $\times$ $\times$ random $\checkmark$ $\times$ $\times$
GPF [182] universal feature/spectral space input $\times$ $\times$ random $\checkmark$ $\checkmark$ $\checkmark$
IGAP [115] signal $\times$ $\times$ random $\checkmark$ $\times$ $\checkmark$
@@
SGL [90] dual-template: $\text{CL}(\mathbf{h}_{u},\mathbf{h}_{v}),\text{GL}(\mathbf{x}_{v},\mathbf{\tilde{x}}_{v})$$\times$ $\checkmark$ $\checkmark$ random $\times$ $\times$ $\checkmark$
@@
HGPrompt [83] dual-template: $\text{sim}(\mathbf{s}_{u},\mathbf{s}_{v})$, graph template readout $\times$ $\times$ random $\checkmark$ $\checkmark$ $\checkmark$
SAP [85] view similarity: $\text{sim}(\text{MLP}(X),\text{GNN}(X,A))$$\times$ $\checkmark$ $\checkmark$ random $\checkmark$ $\times$ $\checkmark$
ULTRA-DP [86] node-node/group similarity: $\text{sim}(\mathbf{h}_{u},\mathbf{h}_{v})$ input $\checkmark$ $\checkmark$ random $\checkmark$ $\checkmark$ $\times$
@@
VNT [87] node attribute reconstruction: $\text{MSE}(\mathbf{x}_{v},\mathbf{\tilde{x}}_{v})$ structure recovery: $\text{MSE}(\{\mathbf{h}_{u},\mathbf{h}_{v}\})$ input $\times$ $\times$ meta-trained $\checkmark$ $\times$ $\times$
@@
ProG [49] subgraph classification: $\text{CLS}(\mathbf{s})$$\times$ $\checkmark$ $\times$ meta-trained $\checkmark$ $\checkmark$ $\checkmark$
DyGPrompt [183] temporal node similarity: $\text{sim}(\mathbf{h}_{t,u},\mathbf{h}_{t,v})$ input $\times$ $\times$ conditional $\checkmark$ $\checkmark$ $\times$
TIGPrompt [184] input $\times$ $\times$ time-based $\checkmark$ $\checkmark$ $\times$
TABLE IX: Summary of prompt tuning on text-attributed graphs.
Paper Instruction Learnable Downstream Task
Text Graph prompt Node Edge Graph
G2P2 [44] $\checkmark$ $\times$ vector $\checkmark$ $\times$ $\times$
G2P2* [82] $\checkmark$ $\times$ condition-net $\checkmark$ $\times$ $\times$
GraphGPT [170] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$
InstructGLM [175] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$
GIMLET [185] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
OFA [45] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
HiGPT [186] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$
V-C1 Prompt Tuning on Graphs
@@
Originating in the field of natural language processing (NLP), prompt-based learning has demonstrated its efficacy in adapting pre-trained language models to diverse language tasks [177]. In essence, a prompt vector is introduced to modify or reformulate the original input for the pre-trained model, drawing the downstream task closer to the pretext tasks under a unified template. Notably, the prompt vector has a negligible size relative to the pre-trained parameters. During the adaptation process, only a very small prompt vector is tuned without updating the pre-trained model, making the approach parameter efficient. Recently, prompt tuning have been extended to graph learning [48, 19]. Building on a unified template that aligns the pretext and downstream losses, these methods design prompts to further narrow the gap between pretext and downstream tasks, as illustrated in Fig. 6(c), right. Specific approaches vary in the choice of the template and prompt design, as summarized in Tables VIIIIX.
@@
Prompt template. Similarity calculation has been widely adopted [48, 19, 80, 179, 84, 86, 83, 183, 181, 180] to unify three major tasks: node classification (NC), link prediction (LP), and graph classification (GC). For instance, GraphPrompt [19] reformulates LP as calculating the similarity between the ego subgraphs of two nodes. Then, it applies the same template to NC and GC, treating them as calculating the similarity between the subgraph of the target instance (using an ego network for NC, or the target graph itself for GC) and the prototypical subgraph for each class. While GraphPrompt only adopts LP as the pretext task, GraphPrompt+ [80] illustrates that other prevailing contrastive pretext tasks can also be redefined using subgraph similarity. Likewise, SGL [90], SAP [85] and VPGNN [178] also employ a form of similarity or matching score to unify pretext and downstream tasks. In particular, SGL further incorporates a generative pretext task based on mask node feature reconstruction, which is employed as an additional template for downstream prompt learning. In contrast, GPF [182] and IGAP [115] do not employ an explicit template to align task formats. Instead, GPF assumes a universal feature space as an implicit template to unify different tasks. Meanwhile, IGAP posits a universal graph signal and spectral space as an implicit template.
@@
Prompt design can be further categorized into feature prompt and structure prompt. On one hand, similar to prompts in NLP [177], feature prompts modify node features, reformulating the original downstream tasks to narrow the gap from the pre-trained model. Some studies incorporate feature prompts in the input layer [48, 182, 86] to explicit modify the input in a task-specific manner, while others apply them to the readout layer to guide subgraph readout [19, 179, 83, 180] or all layers of the pre-trained model [80, 84]. Moreover, from a graph spectral perspective, feature prompts can be applied to graph signals to reduce the spectral gap between various graphs [115]. On the other hand, structure prompts act as a virtual graph structure to reformulate the implicit structural input for downstream tasks. Structure prompts can take the form of virtual nodes or graphs representing class prototypes for classification tasks [90, 85, 86, 178], or virtual nodes that augment link prediction [86]. ProNoG [180] generates a series of prompts for each node using a condition-net, aiming to adapt to node-specific structural patterns on non-homophilic graphs. Furthermore, some studies propose utilizing multiple pretext tasks [90, 85, 86] to acquire comprehensive prior knowledge from diverse perspectives. However, directly combining various pretext tasks may lead to task interference [84], compromising the complementary benefits of employing multiple tasks. To mitigate this issue, MultiGPrompt [84] introduces a composed prompt, which is composed of a series of pretext tokens used during pre-training to promote cooperation among the multiple pretext tasks. Additionally, to leverage prior knowledge across graphs from multiple domains, MDGPT [181] employs a mixing prompt that aggregates a set of domain tokens to bridge the semantic gap across different domains.
Graph types. Prompt-based learning has also been considered on various types of graph. For heterogeneous graphs, HGPrompt [83] extends GraphPrompt, utilizing an additional graph template to convert a heterogeneous graph into homogeneous ones, bridging the gap between the two kinds of graph. Additionally, HetGPT [116] modifies each type of nodes with a type-specific prompt. For dynamic graphs, TIGPrompt [184] proposes node-specific prompts based on timestamps, while DyGPrompt [183] generates conditional prompts for each node and timestamp via dual condition-nets to capture the mutual influence between node and time features. For text-attributed graphs, pre-training often utilizes pre-trained language models. Consequently, in the downstream adaptation stage, various prompt-based learning methods have been specifically designed for these language models, as summarized in Table IX. For instance, G2P2 [44] appends discrete language prompts to textual class labels in zero-shot node classification and tunes learnable prompt vectors in few-shot settings. Moreover, inspired by CoCoOp [187], G2P2* [82] utilizes node embeddings as conditions to train a learnable conditional prompt for the language model. Other works [170, 175, 185, 186] generally describe graph structures using textual instructions for LLMs. Conversely, OFA [45] introduces prompt nodes to unify various downstream tasks by fusing the node-of-interest subgraph, prompt node, and class nodes into a single graph.
V-C2 Parameter-Efficient Fine-Tuning (PEFT)